which of the following is in between 2/5 and 3/5?
A)1/4
B)3/7
C)3/8
D)5/8

Found 5 solutions by Boreal, addingup, JuanKoss, Alan3354, JB567:
Answer by Boreal(15235)

want something between 0.4 and 0.6
1/4 isn't
3/7 is 0.429. That fits
3/8 is 0.375
5/8 is 0.625
B.

Answer by addingup(3677)
Answer by JuanKoss(3)

Can you prompt me? I want, but cant solve it!

Answer by Alan3354(69443)

which of the following is in between 2/5 and 3/5?
--> between 0.4 & 0.6
A)1/4 = 0.25 NG
B)3/7 = 0.42857...
C)3/8 = 0.375 NG
D)5/8 = 0.625 NG
=============================

Answer by JB567(2)

A)1/4
B)3/7
C)3/8
D)5/8